OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (3, 1) and a directrix of y = 5

OpenStudy (anonymous):

which way does it open?

OpenStudy (helder_edwin):

u have the point Q(3,1) and the line L : y-5=0 use the distance formulas for P(x,y) such that \[\large d(P,Q)=d(P,L) \]

OpenStudy (anonymous):

f(x) = one eighth (x − 3)2 + 3 f(x) = −one eighth (x − 3)2 + 3 f(x) = one eighth (x − 3)2 − 3 f(x) = −one eighth (x − 3)2 −

OpenStudy (helder_edwin):

u should know this: if P(a,b) and Q(c,d) then \[\large d(P,Q)=\sqrt{(a-c)^2+(b-d)^2} \] and if L is the line \(y=Ax+B\) then \[\large d(P,L)=\frac{|Aa-b+B|}{\sqrt{A^2+1}} \]

OpenStudy (whpalmer4):

Let's not make this needlessly complicated. Vertex form for a quadratic is \[y=a(x-h)^2+k\] where the vertex is at \((h,k)\) and \(a\) is a constant. We know that the parabola is the set of all points equidistant from the focus and the directrix. That means the vertex must be on a perpendicular line drawn between the focus and the directrix.  Apologies for the shaky parabola! If the vertex is on the line x = 3, and halfway between (3,1) and (3,5), where are its coordinates?

OpenStudy (whpalmer4):

Plug those coordinates into \[y = a(x-h)^2+k\]Compare with your answer choices.

OpenStudy (whpalmer4):

Note that \(a\) controls the vertical scale of the parabola, and whether it opens upward (\(a>0\)) or downward (\(a<0\))